A pizza is cut into eight equal​ pieces, the area of each being 82 cm squared. What was the diameter of the original​ pizza?
Answer by Alan3354(69443)

A pizza is cut into eight equal​ pieces, the area of each being 82 cm squared. What was the diameter of the original​ pizza?
-----------
Area = pi*r^2 = 8*82 sq cms
d = 2r